OpenStudy (anonymous):

Find the first 5 terms of each sequence. An=1/2an-1, where a1= 20

OpenStudy (anonymous):

Do you do connections academy?

OpenStudy (the_fizicx99):

Hmm, what's the difference or ratio? I'm not familiar with what an-1 is

OpenStudy (anonymous):

yes, I do. & That's all the problem gave me. I have no clue how to do it.

OpenStudy (jdoe0001):

\(\bf \Large a_{\color{red}{ n}}=\cfrac{1}{2}a_{{\color{red}{ n}}-1}\qquad \qquad a_1=20\)

OpenStudy (jdoe0001):

\(\Large \bf a_{\color{red}{ 5}}=\cfrac{1}{2}a_{{\color{red}{ 5}}-1}\implies a_{\color{red}{ 5}}=\cfrac{1}{2}a_{{\color{red}{ 4}}}\)

OpenStudy (the_fizicx99):

Ohh, thats the formula they want you to use, plug in 5 for n, a_5 = 1/2*5-1 = a_5 = 1/2*a4

OpenStudy (anonymous):

I do connections as well. I remember this question but not the answer sorry :(

OpenStudy (jdoe0001):

the next term is dependent on the previous term so you'd need to find the 4th to get the 5th and the 3rd to get the 4th and so on your starting point is of course \(\bf a_1=20\)

OpenStudy (jdoe0001):

well.... actually no quite... you don't have to...

OpenStudy (jdoe0001):

\(\bf \Large a_{\color{red}{ n}}=a_1\cdot r^{{\color{red}{ n}}-1}\qquad r=\textit{common ratio}=\cfrac{1}{2}\)

OpenStudy (jdoe0001):

recall -> http://www.mathsisfun.com/algebra/sequences-sums-geometric.html the geometric sequence equation to find the "nth" term

OpenStudy (anonymous):

so then The 5 terms would be?